OpenStudy (anonymous):

Lyle has a ten-cup box of cereal. If one serving is three-fourths cup, how many servings are in the box? thirteen and one-third 7 and one-half twelve and two-thirds ten and three-fourths

OpenStudy (anonymous):

@mkerr017 @haiimawesome

OpenStudy (anonymous):

@haiimawesome

OpenStudy (anonymous):

7 and one-half

OpenStudy (anonymous):

\[\frac{ 10 }{ 1 }\div \frac{ 3 }{ 4}\]Invert and multiply.\[\frac{ 10 }{ 1 }\times \frac{ 4 }{ 3 }=\frac{ 40 }{ 3 }\]\[\frac{ 40 }{ 3}= 13 \frac{ 1 }{ 3}\]